OpenStudy (anonymous):

Simplify: 3/x + 2/x+2 over 3/x+2-2/x a. -11/4 b. -1 c. 5x+6/x-4 d. 5x+2/x+2

OpenStudy (anonymous):

@eliassaab

OpenStudy (ankit042):

First lets solve numerator 3/x +2/(x+2) 3(x+2) +2x/x(x+2) = (5x+6)/x(x+2) Now denominator 3/(x+2) - 2/x (3x-2(x+2))/ x(x+2) = (x-4)/x(x+2) so final answer 5x+6/x-4